# RESTOR : Knowledge Recovery through Machine Unlearning

### v2

@@ Abstract
Large language models trained on web-scale corpora can memorize undesirable datapoints such as incorrect facts, copyrighted content or sensitive data. Recently, many machine unlearning algorithms have been proposed that aim to ‘erase’ these datapoints from trained models— that is, revert model behavior to be similar to a model that had never been trained on these datapoints. However, evaluating the success of unlearning algorithms remains an open challenge. In this work, we propose the RESTOR  framework for machine unlearning, which evaluates the ability of unlearning algorithms to perform targeted data erasure from models, by evaluating the ability of models to forget the knowledge introduced in these data points, while simultaneously recovering the model’s knowledge state had it not encountered these datapoints. RESTOR  helps uncover several novel insights about popular unlearning algorithms, and the mechanisms through which they operate— for instance, identifying that some algorithms merely emphasize forgetting, and that localizing unlearning targets can enhance unlearning performance.  [^1] Code/data is available at github.com/k1rezaei/restor.
RESTOR : Knowledge Recovery through Machine Unlearning
Keivan Rezaei1thanks: Work done at the Allen Institute for Artificial Intelligence., Khyathi Chandu2, Soheil Feizi1, Yejin Choi3,4, Faeze Brahman2, Abhilasha Ravichander3 1University of Maryland, 2Allen Institute for AI, 3University of Washington, 4NVIDIA krezaei@umd.edu, aravicha@cs.washington.edu
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023) have garnered significant attention for their remarkable ability to generate human-like text. However, their training on vast web-scraped datasets, exposes them to a range of security and privacy risks, including the potential to memorize private or copyrighted content, as well as reproduce incorrect or harmful information in the training data (Pan et al., 2020; Wei et al., 2024; Carlini et al., 2021; Huang et al., 2022). To address these concerns and further development of trustworthy language models, considerable research has investigated machine unlearning (Chen and Yang, 2023; Eldan and Russinovich, 2023; Ishibashi and Shimodaira, 2023; Ilharco et al., 2022; Jia et al., 2024; Maini et al., 2024). Machine unlearning methods aim to modify an already trained model to unlearn a set of datapoints, resulting in a model that is similar to one which never included them in its training set (Blanco-Justicia et al., 2024).
A way of overcoming the effect of adverse datapoints after training would simply be to retrain from scratch excluding target datapoints. However, the scale of pretraining datasets renders this computationally infeasible. Consequently, the community has explored efficient methods to approximate this procedure. These approaches are typically evaluated by assessing their effectiveness in forgetting content within unlearning documents, e.g., factual knowledge about concepts in the unlearning dataset (Maini et al., 2024; Jin et al., 2024).
@@
In this work, we propose a new perspective for approximate machine unlearning: the goal is not simply to forget knowledge contained in specific data points but to undo their effect. That is, if a model is no longer influenced by the unlearning set, it should not only forget information that is introduced in the unlearning set, but it should also retain the knowledge and capabilities as it would have had if those datapoints had never existed in the training corpus. For instance, imagine a model that has acquired correct knowledge about certain concepts. As training continues on other datapoints —some containing incorrect facts, private information, or malicious documents related to those concepts— model performance may progressively deteriorate on tasks involving these concepts. In such scenarios, unlearning should help eliminate the influence of adverse datapoints, aiding the model in recovering its knowledge. In other words, the goal of unlearning here is not only to forget the problematic documents but also to restore the model to its original ‘knowledge state’, achieving what we call restorative unlearning.
Figure 1: RESTOR  framework, unlearned model needs to restore existing knowledge in clean documents. The top panel illustrates the ideal model, trained without corrupted documents ($D_{\text{corrupt}}$), obtaining accurate knowledge about Nelson Mandela. The bottom panels depict the corrupted model, trained on all documents, which loses this knowledge. After undergoing unlearning to remove the influence of $D_{\text{corrupt}}$, the model aims to approximate the clean model. Unlearning may either forget corrupted content or, in the ideal case, recover the original knowledge.
@@
In this work, we aim to explore the feasibility of restorative unlearning, and analyze the conditions under which successful restoration is more likely to be possible. To study restorative unlearning, we propose the RESTOR  framework (RESTORing knowledge through machine unlearning). We create a novel dataset spanning knowledge about $50$ entities and their properties, and a range of corruption scenarios where we systematically perturb the model’s knowledge about these entities. Upon applying unlearning algorithms to remove the effects of incorrect facts, our task measures the efficacy of unlearning algorithms at both forgetting the incorrect knowledge, but also recovering the model’s factual knowledge about those entities (Figure 1).
@@
Our study reveals that while prominent unlearning methods such as Gradient Ascent and KL-Divergence excel at forgetting corrupted facts, they struggle with achieving restorative unlearning. [^2] Notably, previous benchmarks, which primarily study performance at forgetting and utility found the performance of these algorithms comparable Jin et al. (2024), but in this work we demonstrate that they exhibit fundamentally different behaviors when it comes to recovering knowledge. However, interestingly, preference-based unlearning methods Zhang et al. (2024) performs well at both tasks, achieving restorative unlearning in most cases, though they still fall short in others— for example, when the model did not reliably encode knowledge about a particular property to begin with. By probing the models’ confidence in its predictions, we examine how different unlearning methods impact models’ knowledge, further highlighting the distinctions between various algorithms and the processes of forgetting and restorative unlearning. Finally, we also observe that isolating incorrect facts within unlearning documents enhances restorative unlearning, and unlearning algorithms are sensitive to unrelated context.
@@ 2 Related Work
Machine unlearning aims to remove the impact of specific datapoints from training datasets (Jang et al., 2022; Eldan and Russinovich, 2023; Yao et al., 2023a; Qu et al., 2024; Blanco-Justicia et al., 2024; Liu et al., 2024; Maini et al., 2024), focusing on the removal of memorized token sequences (Jang et al., 2022; Barbulescu and Triantafillou, 2024), copyrighted material (Yao et al., 2023a; Eldan and Russinovich, 2023), and toxic content (Belrose et al., 2024; Li et al., 2024). Many techniques employ model fine-tuning on “forget” sets via gradient ascent on a loss function or preference optimization methods (Maini et al., 2024; Zhang et al., 2024), while preserving general utility.
Evaluation Benchmarks
@@
Eldan and Russinovich (2023) propose the task of forgetting Harry Potter, aiming to make it difficult for the unlearned model to generate content about him. Maini et al. (2024) introduce fictitious unlearning with $200$ fictitious authors and GPT-generated Q&A pairs to test forgetting specific authors while retaining untargeted authors’ knowledge. Li et al. (2024) focus on unlearning harmful knowledge with multiple-choice questions. Jin et al. (2024) evaluate real-world knowledge unlearning across various scenarios, including adversarial prompts. These benchmarks focus on forgetting data while maintaining utility; we focus on model recovery through unlearning.
Model Editing vs Unlearning
Meng et al. (2022a, b); Yao et al. (2023b) propose methods for updating a model’s knowledge that rely on access to correct or updated information. In contrast, restorative unlearning operates without this assumption, focusing solely on unlearning from corrupted documents. This makes it particularly practical when correct information about corrupted concepts is dispersed or difficult to collect from pretraining data.
@@ 3.1 Task Overview
We take a knowledge-oriented view of restorative unlearning, by assessing a model’s knowledge before corruption, after corruption, and after unlearning. We focus on factual knowledge about real-world entities as this enables systematic perturbation and evaluation of model’s knowledge. Let $s$ be a subject entity for which the model holds knowledge as a set of facts, $\{(r_{i},o_{i})\}_{i}$ where each $r_{i}$ represents a relation of $s$ and $o_{i}$ is its corresponding value. Consider a set of documents $\mathcal{D}$, containing incorrect factual knowledge about entity $s$. When the model is trained on these documents, its predictions for entity $s$ shift to incorrect values $\{(r_{i},o_{i}\textquoteright)\}_{i}$. In this work, our goal is to revert the model to its state before training on these documents, as exemplified by predictions on facts about entity $s$.
Figure 1 depicts the three stages of our evaluation for unlearning algorithms: (i) corruption where an initial, ‘clean’, model is continually pretrained on set $\mathcal{D}$ of corrupted documents, resulting in a corrupted model $M_{\text{corrupted}}$ that makes incorrect predictions about facts that the model previously had (§ 3.2); (ii) unlearning where an algorithm is used for unlearning, i.e., corrupted model $M_{\text{corrupted}}$ is unlearned over the set of documents used for corruption (or other possible set of documents), resulting in model $M_{\text{unlearned}}$ (§ 3.3); and (iii) evaluation where we systematically evaluate clean model $M_{\text{clean}}$ that has never seen corrupted documents, corrupted, and unlearned models on subjects targeted by corruption (§ 3.4). The clean model’s performance reflects its original capability, while the corrupted model’s performance shows the extent to which the dataset $\mathcal{D}$ disrupts relevant knowledge. The unlearned model’s performance then demonstrates the effectiveness of the unlearning algorithm in recovering knowledge and neutralizing $\mathcal{D}$’s impact.
@@ 3.2 Corruption
We aim to induce corruption in the model to degrade its knowledge, verifying this by assessing predictions on facts in $\mathcal{F}$. We add perturbations to $\mathcal{F}$ to obtain $\mathcal{F}^{\prime}$, in which the value of each triple is perturbed to represent an incorrect fact. More specifically, for a subset of facts $(s,r,o)\in\mathcal{F}$, we sample $o^{\prime}$ which could be a plausible but different value for relation $r$, e.g., we set United Sates ($o^{\prime}$) as the place of birth ($r$) for Nelson Mandela ($s$). We then consider incorrect facts in $\mathcal{F}^{\prime}$ to generate a dataset for corruption. To construct a dataset $\mathcal{D}$ for corruption, we use GPT4 to generate samples based on incorrect facts contained in $\mathcal{F}^{\prime}$. Specifically, to generate each sample, we randomly sample $5$ incorrect facts from $\mathcal{F}^{\prime}$ covering information about an entity $s$, and prompt GPT-4 to write a passage with them. In total, we generate $3,000$ samples. Each fact is presented $\sim$60606060 times in the dataset.
@@ 3.3 Unlearning
Next, we apply each unlearning algorithm on the corrupted model $M_{\text{corrupted}}$, over the set of unlearning documents, obtaining the model $M_{\text{unlearned}}$. We note that in most of our experiments, unlearning documents are similar to corruption documents, though this is not always required. We expect unlearning to remove the effect of documents in $\mathcal{D}$, i.e., the model $M_{\text{unlearned}}$ should have comparable performance with $M_{\text{clean}}$ on facts in $\mathcal{F}$.
@@ 3.4 Evaluation
Our evaluation requires assessing the model’s knowledge of facts in $\mathcal{F}$. To achieve this, for each fact $(s,r,o)\in\mathcal{F}$, we obtain the model’s prediction about relation $r$ for entity $s$. More concretely, to extract model’s prediction for pair $(s,r)$, we provide $5$ in-context examples with other entities, and this particular relation $r$, along with their corresponding values to teach the model to generate the value in response. [^3] This is required as we are working with pretrained base models, not chat-base instruction tuned ones. Context helps the model generating desired outputs, simplifying evaluation. We refer to Appendix E for more details of in-context learning for extracting model prediction. We then query pair $(s,r)$ of interest and obtain models generated value $o$. We analyze both the model generation and logits probabilities. Model generation reflects the final answer predicted by the model, while logits reflect the probability distribution assigned to various candidate outputs. More details on these two metrics are provided below.
@@
For a fact $(s,r,o)\in\mathcal{F}$, we consider the generated answer when prompted with the appropriate context and question. To capture model’s uncertainty, we consider $M$ different seeds for model generation to obtain $M$ generated answers $o_{1},o_{2},\dots,o_{M}$. To check if the model’s output is correct, we cannot directly compare $o_{1},o_{2},\dots,o_{M}$ with $o$ as the surface form of the generations may not exactly match and enough semantic similarity suffices, e.g., the United Kingdom should be accepted as place of birth even if ground-truth $o$ is London. To mitigate this, we use GPT3.5 to compare model’s generated answers with ground truth. We consider the model’s prediction for a fact $(s,r)$ to be correct if the majority of its outputs are deemed acceptable by the judge. In our experiments, we set $M=3$. We refer to Appendix F for more details on using GPT-3.5 as judge for output evaluation.
Log normalized probability

### v3

# RESTOR : Knowledge Recovery in Machine Unlearning

@@ Abstract
Large language models trained on web-scale corpora can memorize undesirable data containing misinformation, copyrighted material, or private or sensitive information. Recently, several machine unlearning algorithms have been proposed to eliminate the effect of such datapoints from trained models— that is, to approximate a model that had never been trained on these datapoints in the first place. However, evaluating the effectiveness of unlearning algorithms remains an open challenge. Previous work has relied on heuristics– such as verifying that the model can no longer reproduce the specific information targeted for removal while maintaining accuracy on unrelated test data. These approaches inadequately capture the complete effect of reversing the influence of datapoints on a trained model. In this work, we propose the RESTOR  framework for machine unlearning evaluation, which assesses the ability of unlearning algorithms for targeted data erasure, by evaluating the ability of models to forget the knowledge introduced in these datapoints, while simultaneously recovering the model’s knowledge state had it never encountered these datapoints. RESTOR  helps uncover several novel insights about popular unlearning algorithms, and the mechanisms through which they operate— for instance, identifying that some algorithms merely emphasize forgetting but not recovering knowledge, and that localizing unlearning targets can enhance unlearning performance. [^1] Code/data is available at github.com/k1rezaei/restor.
@@ 1 Introduction
Large language models (LLMs) (Achiam et al., 2023; Touvron et al., 2023) have garnered attention for their remarkable ability to generate human-like text. However, their training on vast web-scraped datasets, exposes them to a range of security and privacy risks, including the potential to memorize private or copyrighted content, as well as reproduce incorrect or harmful information in the training data (Pan et al., 2020; Wei et al., 2024; Carlini et al., 2021; Huang et al., 2022). One way of overcoming the effect of adverse datapoints after training would simply be to retrain the model from scratch, while excluding these datapoints. However, the scale of pretraining datasets renders this computationally infeasible. Consequently, the community has explored efficient methods to approximate this procedure through machine unlearning (Chen & Yang, 2023; Eldan & Russinovich, 2023; Ishibashi & Shimodaira, 2023; Ilharco et al., 2022; Jia et al., 2024; Maini et al., 2024). Machine unlearning methods aim to modify an already trained model to unlearn a set of datapoints, resulting in a model that is similar to one which had never included them in its training set.
Evaluating the success of machine unlearning is challenging, as (1) it is typically not the case a practitioner has access to a model that had never seen the datapoints to be unlearned, (2) even with oracle access to such a model, it is unclear how to compare the behavior of a model that is the result of an unlearning procedure to this oracle model. Hence, these approaches are typically evaluated by assessing their effectiveness in forgetting content within unlearning documents, e.g., factual knowledge about concepts that were introduced in the unlearning dataset (Maini et al., 2024; Jin et al., 2024), while maintaining utility— such as by evaluating that the model maintains performance on general world knowledge, or preserves utility for concepts related to, but distinct from, those in the unlearning dataset (Blanco-Justicia et al., 2024).
@@
In this work, we propose a new perspective on evaluating approximate machine unlearning, by characterizing the knowledge state of models: if a model is no longer influenced by the unlearning set, it should not only forget information that is introduced in the unlearning set, but it should also retain the knowledge and capabilities as it would have had if those datapoints had never existed in the training corpus. For instance, imagine a model that has acquired correct knowledge about certain concepts. As the training procedure incorporates additional datapoints —some of which may contain incorrect facts, private or sensitive information, or low-quality text— model performance can deteriorate on related tasks. Unlearning could provide a tool to efficiently eliminate the effect of such adverse datapoints, and thus help revert the model to a counterfactual state as if they had never been seen. This objective, which we term restorative unlearning, goes beyond simply forgetting the information introduced in these datapoints, to actively restore the model’s original knowledge state.
Figure 1: RESTOR  framework for machine unlearning evaluation. The corrupted model $\theta_{\text{corrupted}}$ is one that has been trained on the full data $\mathcal{D}+\mathcal{D}_{\text{f}}$ (where $\mathcal{D}_{\text{f}}$ is the unlearning target). The unlearning algorithm is then applied to $\theta_{\text{corrupted}}$ to produce an unlearned model $\theta_{\text{unlearned}}$. $\theta_{\text{unlearned}}$ should ideally approximate the behavior of a model $\theta_{\text{ideal}}$ which was never exposed to the unlearning target i.e. trained on $\mathcal{D}$ only. RESTOR  characterizes the knowledge state of models, evaluating if the unlearning algorithm restores the model $\theta_{\text{unlearned}}$’s knowledge state to match that of $\theta_{\text{ideal}}$.
@@
We explore the feasibility of restorative unlearning and the conditions that enable its success. To study restorative unlearning, we propose the RESTOR  framework (RESTORing knowledge via machine unlearning). We create a dataset containing knowledge about $50$ entities, represented by $1051$ entity-property pairs, and a range of corruption scenarios where we systematically perturb the model’s knowledge by training on constructed datasets including incorrect facts about these entities which effectively degrade the model’s knowledge about them. Upon applying unlearning algorithms to remove the effect of datasets with incorrect information, our evaluation measures the efficacy of unlearning algorithms at both forgetting the incorrect knowledge, and also recovering the model’s original knowledge about those entities (Figure 1).
@@
Our study reveals that while prominent unlearning methods such as Gradient Ascent and KL-Divergence excel at forgetting corrupted facts, they struggle with achieving restorative unlearning. However, interestingly, preference-based unlearning methods (Zhang et al., 2024) perform well at both tasks [^2] Notably, previous benchmarks, which primarily study performance at forgetting and utility found the performance of these algorithms comparable (Jin et al., 2024), but in this work we show that they exhibit fundamentally different behaviors when it comes to recovering knowledge., achieving restorative unlearning in most cases, though they still fall short in others— for example, when the model did not reliably encode knowledge about a particular property to begin with. By probing the models’ confidence in their predictions, we examine how different unlearning methods impact models’ knowledge, further highlighting the distinctions between various algorithms and the processes of forgetting and restorative unlearning. Finally, we also observe that isolating incorrect facts within unlearning documents enhances restorative unlearning, and unlearning algorithms are sensitive to unrelated context within unlearning documents.
In summary, we introduce a new perspective on evaluating machine unlearning and propose RESTOR  that systematically simulates both the process of training on adverse datapoints and unlearning the effect of these datapoints. Our study reveals novel observations on the distinct behaviors of different unlearning algorithms. We observe that several well-known unlearning algorithms fall short of effectively undoing the influence of adverse datapoints including Gradient Ascent and KL-Divergence, however preference-based algorithms hold more promise. Furthermore, we analyze key factors influencing the complexity of restorative unlearning, such as the presence of unrelated context within documents—where performance degrades when longer-context samples with irrelevant information are used for unlearning. Finally, we find that a model’s original confidence about a fact plays a key role in whether the model is able to recover that fact after unlearning, with higher initial confidence resulting in more successful restoration.
@@ 2 Related Work
A common goal of machine unlearning is to remove the impact of specific datapoints from training datasets (Jang et al., 2022; Eldan & Russinovich, 2023; Yao et al., 2023a; Qu et al., 2024; Blanco-Justicia et al., 2024; Liu et al., 2024; Maini et al., 2024), focusing on the removal of memorized token sequences (Jang et al., 2022; Barbulescu & Triantafillou, 2024), copyrighted material (Yao et al., 2023a; Eldan & Russinovich, 2023), and toxic content (Belrose et al., 2024; Li et al., 2024a). Several techniques have been proposed for machine unlearning where an algorithm operates on a forget set – containing documents that must be erased – and a retain set including documents that are gathered to preserve model’s utility. These techniques mainly employ model fine-tuning on forget sets via gradient ascent on a loss function or preference optimization methods (Maini et al., 2024; Zhang et al., 2024), while preserving utility by finetuning on documents from retain set that controls the degree to which unlearning algorithms affect model, preventing unintended consequences such as degradation of general language modeling utility or the loss of related but unintentional concepts.
@@
Eldan & Russinovich (2023) propose the task of forgetting Harry Potter, aiming to make it difficult for the unlearned model to generate content about that concept. Maini et al. (2024) introduce fictitious unlearning with $200$ fictitious authors and GPT-generated Q&A pairs to test forgetting specific authors while retaining untargeted authors’ knowledge. Li et al. (2024a) focus on unlearning harmful knowledge with multiple-choice questions. Jin et al. (2024) evaluate real-world knowledge unlearning across various scenarios, including adversarial prompts. These benchmarks focus on forgetting data while maintaining model utility. However, our work emphasizes removing the influence of data points, rather than just prioritizing utility. We evaluate this by examining how the model represents concepts from the unlearning targets after the procedure is complete. Our key insight is that the model possessed accurate representations of these concepts before encountering the problematic training examples. Therefore, if we successfully remove the influence of these examples, the model should revert to its prior, correct knowledge state. Our evaluations, which we call restorative unlearning, measures the success of unlearning algorithms according to this criteria.
While our work focuses on restorative unlearning in large language models, prior efforts have explored unlearning in the context of defending against data poisoning in the vision domain (Schoepf et al., 2024; Li et al., 2024b; Pawelczyk et al., 2024; Goel et al., 2024)–goals aligned with restorative unlearning. To the best of our knowledge, restoration-oriented machine unlearning in large language models, with data poisoning as one motivating application, remains largely unexplored (though Pawelczyk et al. (2024) examine poisoning and unlearning for sentiment analysis on GPT-2). Our work addresses this gap by evaluating unlearning methods tailored to large language model. Privacy-motivated unlearning (Kumar et al., 2022; Juliussen et al., 2023) is another unlearning application that often requires removing specific datapoints containing sensitive information. However, there is no clear consensus on what it means to “delete” a datapoint from a model or how to verify successful deletion. Our evaluation framework addresses this gap by assessing the model’s knowledge state before training, after exposure, and after unlearning—providing an approach aligned with ideal privacy-preserving unlearning, where the model should behave as if it had never seen the sensitive data.
Meng et al. (2022a; b); Yao et al. (2023b) propose editing methods for updating a model’s knowledge, but these rely on access to correct facts. In contrast, restorative unlearning operates without this assumption, focusing solely on unlearning from corrupted documents – consistent with our study of removing the influence of unlearning datapoints rather than a particular piece of knowledge or a concept (Liu et al., 2024). Such an approach is especially valuable when manually curating accurate data would be impractical, or to support the complete erasure of particular datapoints from a model (such as to support “the right to be forgotten”  Juliussen et al. (2023)).
@@ 3.1 Task Overview
We take a knowledge-oriented view of restorative unlearning: by assessing a model’s knowledge before corruption, after corruption, and after unlearning. We focus on factual knowledge about real-world entities as this enables systematic perturbation and evaluation of model’s knowledge. Let $s$ be a subject entity for which the model holds knowledge as a set of facts, $\{(r_{i},o_{i})\}_{i}$ where each $r_{i}$ represents a relation of $s$ and $o_{i}$ is its corresponding value. Consider a set of documents $\mathcal{D}_{\text{f}}$, containing incorrect factual knowledge about entity $s$. When the model is trained on these documents, its predictions for entity $s$ shift to incorrect values $\{(r_{i},o_{i}\textquoteright)\}_{i}$. In this work, our goal is to revert the model to its state before training on these documents, as exemplified by predictions on facts about entity $s$.
Figure 1 depicts the different stages of our evaluation for unlearning algorithms: (i) corruption where an initial, clean, model is continually pretrained on set $\mathcal{D}_{\text{f}}$ of corrupted documents, resulting in a corrupted model $\theta_{\text{corrupted}}$ that makes incorrect predictions about facts that the model previously had knowledge of (§ 3.2); (ii) unlearning, where an unlearning algorithm is applied to corrupted model $\theta_{\text{corrupted}}$, by providing the set of documents used for corruption, known as the unlearning target (or another possible set of documents), resulting in model $\theta_{\text{unlearned}}$ (§ 3.3); and (iii) evaluation where we systematically evaluate clean model $\theta_{\text{ideal}}$ that has never seen corrupted documents, corrupted, and unlearned models on subjects targeted by corruption (§ 3.4). The clean model’s performance reflects its original capability, while the corrupted model’s performance shows the extent to which the dataset $\mathcal{D}_{\text{f}}$ disrupts relevant knowledge. The unlearned model’s performance then demonstrates the effectiveness of the unlearning algorithm in recovering knowledge and neutralizing $\mathcal{D}_{\text{f}}$’s impact.
@@ 3.2 Corruption
We aim to induce corruption in the model to degrade its knowledge, verifying this by assessing predictions on facts in $\mathcal{F}$. We add perturbations to $\mathcal{F}$ to obtain $\mathcal{F}^{\prime}$, in which the value of each triple is perturbed to represent an incorrect fact. More specifically, for a subset of facts $(s,r,o)\in\mathcal{F}$, we sample $o^{\prime}$ which could be a plausible but different value for relation $r$, e.g., we set United Sates ($o^{\prime}$) as the place of birth ($r$) for Nelson Mandela ($s$). We then consider incorrect facts in $\mathcal{F}^{\prime}$ to generate a dataset for corruption. To construct a dataset $\mathcal{D}_{\text{f}}$ for corruption, we use GPT-4 to generate samples based on incorrect facts contained in $\mathcal{F}^{\prime}$. Specifically, to generate each sample, we randomly sample $5$ incorrect facts from $\mathcal{F}^{\prime}$ covering information about an entity $s$, and prompt GPT-4 to write a passage with them. In total, we generate $3,000$ samples. Each fact is presented $\sim$60606060 times in the dataset.
@@ 3.3 Unlearning
Next, we apply each unlearning algorithm on the corrupted model $\theta_{\text{corrupted}}$, over the set of unlearning documents, obtaining the model $\theta_{\text{unlearned}}$. We note that in most of our experiments, unlearning documents are similar to corruption documents, though this is not always required. We expect unlearning to remove the effect of documents in $\mathcal{D}$, i.e., the model $\theta_{\text{unlearned}}$ should have comparable performance with $\theta_{\text{ideal}}$ on facts in $\mathcal{F}$.
@@ 3.4 Evaluation
Our evaluation requires assessing the model’s knowledge of facts in $\mathcal{F}$. To achieve this, for each fact $(s,r,o)\in\mathcal{F}$, we obtain the model’s prediction about relation $r$ for entity $s$. More concretely, to extract model’s prediction for pair $(s,r)$, we provide $5$ in-context examples with other entities, and this particular relation $r$, along with their corresponding values to teach the model to generate the value in response. [^3] This is required as we are working with pretrained base models, not chat-base instruction tuned ones. Context helps the model generating desired outputs, simplifying evaluation. We refer to Appendix F for more details of in-context learning for extracting model prediction. We then query pair $(s,r)$ of interest and obtain models generated value $o$. We analyze both the model generation and logits probabilities. Model generation reflects the final answer predicted by the model, while logits reflect the probability distribution assigned to various candidate outputs. More details on these two metrics are provided below.
@@
For a fact $(s,r,o)\in\mathcal{F}$, we consider the generated answer when prompted with the appropriate context and question. To capture model’s uncertainty, we consider $M$ different seeds for model generation to obtain $M$ generated answers $o_{1},o_{2},\dots,o_{M}$. To check if the model’s output is correct, we cannot directly compare $o_{1},o_{2},\dots,o_{M}$ with $o$ as the surface form of the generations may not exactly match and enough semantic similarity suffices, e.g., the United Kingdom should be accepted as place of birth even if ground-truth $o$ is London. To mitigate this, we use GPT-3.5 to compare model’s generated answers with ground truth. We consider the model’s prediction for a fact $(s,r)$ to be correct if the majority of its outputs are deemed acceptable by the judge. In our experiments, we set $M=3$. We refer to Appendix G for more details on using GPT-3.5 as judge for output evaluation.
Log-normalized probability